# Forum for Information Retrieval Evaluation, December 12-15, 2024, India

### v1

@@ 1 Introduction
By focusing on the cryptocurrency domain, this task [5] not only contributes to the broader field of sentiment analysis and text classification but also offers practical implications for stakeholders in the financial sector. Improved classification and information retrieval from social media can lead to better market insights, more effective communication strategies, and enhanced user experiences in digital finance. This study aims to showcase the potential of natural language processing techniques in transforming raw social media data into actionable intelligence in the context of cryptocurrency [6]. In our approach to solving the classification task for Reddit and Twitter posts, we employed a prompt-based technique [7], which has proven to be effective in leveraging the strengths of large language models for text classification. Specifically, we utilized carefully designed prompts to guide the model in understanding the context and nuances of social media posts related to cryptocurrency. This approach enabled us to harness the power of GPT-4-Turbo [8, 9] to classify posts into predefined categories, such as objective, positive, negative, etc.
@@ 2 Related Work
The task of classifying and analyzing social media posts, particularly in specialized domains such as cryptocurrency, has garnered significant attention in recent years [3]. Social media platforms like Twitter and Reddit are rich sources of user-generated content that reflect public sentiment, market trends, and community discussions [11]. However, the unstructured and often noisy nature of social media data presents challenges for traditional text classification and information retrieval methods [12, 13]. Early efforts in text classification on social media data primarily relied on traditional machine learning approaches, such as support vector machines (SVMs) and logistic regression, combined with handcrafted features like term frequency-inverse document frequency (TF-IDF) and n-grams [14, 15, 16, 17]. However, these approaches often struggled with the informal and context-dependent language prevalent on platforms like Twitter and Reddit.
With the advent of deep learning, particularly with the development of convolutional neural networks (CNNs), recurrent neural networks (RNNs), and transformer-based language models (LM), the field saw a shift towards more robust models that could capture the nuances of language more effectively [18, 19, 20, 21, 22, 23, 24, 25]. More recently, transformer-based models such as BERT (Bidirectional Encoder Representations from Transformers) have set new benchmarks in text classification by leveraging contextual embeddings to better understand the intricacies of language [26, 27]. These models have been particularly effective in sentiment analysis, where understanding the context and tone of a post is crucial.
@@
Few-shot learning techniques [10] have gained traction as a means to improve model performance in scenarios where labeled data is scarce. In particular, the use of n-shot learning, where the model is exposed to a small number of labeled examples during training, has shown promise in various NLP tasks [28, 29]. The 64-shot technique we employed is an extension of this concept, allowing the model to learn from a diverse set of examples, thereby improving its generalization to unseen data . Recent studies have demonstrated the effectiveness of few-shot learning in improving the performance of large language models (LLMs) on various tasks, including text classification [30, 31, 32] and question answering [33, 34, 35, 36, 37, 38].
@@ 5.1 Why Prompting?
  • Flexibility and Adaptability: Prompting offers significant flexibility in task formulation [41, 42, 43]. By simply changing the structure or wording of a prompt, we can adapt the model to perform a wide variety of tasks without the need for retraining [44, 45, 46, 47]. This adaptability is crucial when dealing with dynamic content, such as social media posts, new legal judgments, etc where domains, language, and context can rapidly evolve [48, 49, 50, 51].
  • Improved Contextual Understanding: Social media posts, especially those related to cryptocurrency, often contain informal language, abbreviations, slang, and domain-specific jargon [52]. Traditional classification models might struggle with this variability. Prompt-based approaches, however, leverage the model’s deep contextual understanding, allowing it to better interpret and classify such content [7, 53]. For instance, a well-crafted prompt can help the model distinguish between positive and negative sentiment even when the language is complex or non-standard.
  • Enhanced Relevance Determination: Finding relevant answers to specific questions within a corpus of social media posts is a complex task, as it requires understanding the subtle relationships between questions and potential answers [54, 55]. Prompting allows us to frame this task in a way that aligns with the model’s strengths, treating it as a natural language inference problem where the model evaluates the likelihood that a given answer is relevant to the question [7]. Using prompt-based few-shot learning, we can further fine-tune this ability, improving the accuracy of relevance determination in contexts where direct answer-question pairs are not always straightforward [36].

### v2

@@ 1 Introduction
By focusing on the cryptocurrency domain, this task [5] not only contributes to the broader field of sentiment analysis and text classification but also offers practical implications for stakeholders in the financial sector. Improved classification and information retrieval from social media can lead to better market insights, more effective communication strategies, and enhanced user experiences in digital finance. This study aims to showcase the potential of natural language processing techniques in transforming raw social media data into actionable intelligence in the context of cryptocurrency [6]. In our approach to solving the classification task for Reddit and Twitter posts, we employed a prompt-based technique [7], which has proven to be effective in leveraging the strengths of large language models for text classification. Specifically, we utilized carefully designed prompts to guide the model in understanding the context and nuances of social media posts related to cryptocurrency. This approach enabled us to harness the power of GPT-4-Turbo [8] to classify posts into predefined categories, such as objective, positive, negative, etc.
@@ 2 Related Work
The task of classifying and analyzing social media posts, particularly in specialized domains such as cryptocurrency, has garnered significant attention in recent years [3]. Social media platforms like Twitter and Reddit are rich sources of user-generated content that reflect public sentiment, market trends, and community discussions [10]. However, the unstructured and often noisy nature of social media data presents challenges for traditional text classification and information retrieval methods [11]. Early efforts in text classification on social media data primarily relied on traditional machine learning approaches, such as support vector machines (SVMs) and logistic regression, combined with handcrafted features like term frequency-inverse document frequency (TF-IDF) and n-grams [12, 13, 14, 15]. However, these approaches often struggled with the informal and context-dependent language prevalent on platforms like Twitter and Reddit.
With the advent of deep learning, particularly with the development of convolutional neural networks (CNNs) and recurrent neural networks (RNNs), the field saw a shift towards more robust models that could capture the nuances of language more effectively [16, 17, 18, 19, 20]. More recently, transformer-based models like BERT (Bidirectional Encoder Representations from Transformers) have set new benchmarks in text classification by leveraging contextual embeddings to better understand the intricacies of language [21, 22]. These models have been particularly effective in sentiment analysis, where understanding the context and tone of a post is crucial.
@@
Few-shot learning techniques [9] have gained traction as a means to improve model performance in scenarios where labeled data is scarce. In particular, the use of n-shot learning, where the model is exposed to a small number of labeled examples during training, has shown promise in various NLP tasks [23]. The 64-shot technique we employed is an extension of this concept, allowing the model to learn from a diverse set of examples, thereby improving its generalization to unseen data . Recent studies have demonstrated the effectiveness of few-shot learning in improving the performance of large language models (LLMs) on various tasks, including text classification [24, 25, 26] and question answering [27, 28].
@@ 5.1 Why Prompting?
  • Flexibility and Adaptability: Prompting offers significant flexibility in task formulation. By simply changing the structure or wording of a prompt, we can adapt the model to perform a wide variety of tasks without the need for retraining [7]. This adaptability is crucial when dealing with dynamic content, such as social media posts, where topics, language, and context can rapidly evolve.
  • Improved Contextual Understanding: Social media posts, especially those related to cryptocurrency, often contain informal language, abbreviations, slang, and domain-specific jargon [31]. Traditional classification models might struggle with this variability. Prompt-based approaches, however, leverage the model’s deep contextual understanding, allowing it to better interpret and classify such content [7]. For instance, a well-crafted prompt can help the model distinguish between positive and negative sentiment even when the language is complex or non-standard.
  • Enhanced Relevance Determination: Finding relevant answers to specific questions within a corpus of social media posts is a complex task, as it requires understanding the subtle relationships between questions and potential answers [32]. Prompting allows us to frame this task in a way that aligns with the model’s strengths, treating it as a natural language inference problem where the model evaluates the likelihood that a given answer is relevant to the question [7]. Using prompt-based few-shot learning, we can further fine-tune this ability, improving the accuracy of relevance determination in contexts where direct answer-question pairs are not always straightforward [28].